# Quasi Inverse of Qubit Channels for Mixed States

### v1

@@ Abstract
We found the quasi inverse of qubit channels as a unitary map, $\mathcal{E}^{i}$, by minimizing the average trace distance between the input state to the channel and the output of the quasi inverse channel for arbitrary qubit channel $\mathcal{E}$ and for arbitrary input states. The channel $\mathcal{E}$ was assumed completely positive and trace-preserving. To find the quasi inverse for mixed states, we proposed an alternative definition of the quasi inverse based on the mean square of the trace distance (MSTD) of the mixed input state of the original channel and the output of the quasi inverse. The definition based on the trace distance allowed easy generalization of the quasi inverse to mixed input states. The quasi inverse of the Pauli, generalized amplitude damping, mixed unitary, and tetrahedron channels calculated based on trace distance agreed with the one computed using fidelity in the special case of input states being pure.
@@ I Introduction
Qubits are the building blocks of quantum computers, but they are quite vulnerable to unwanted interactions with environment. This interaction leads to the decoherence of the quantum state and makes the computations erraneous on quantum computers. This interaction process can be modeled as a quantum channel [1]. In general, quantum channels are irreversible operations [2, 3, 4]. However, the idea of quasi inversion of the quantum channel was recently introduced [2] to reverse a part of the effect of that channel. The concept of quasi inverse is essential for enhancing the fidelity of quantum gates and reducing errors in quantum computations. This idea is also helpful in quantum inverse problems, such as state estimation, where quasi inverses are used to improve efficiency and resilience [5].
@@
The quasi inverse is defined so as to increase the average fidelity of the channel in Ref. [2]. The derivation of the quasi inverse was based on maximising the average fidelity between the input state of the channel and the output state of the quasi inverse of that channel. However, the derivation was restricted to the case when the input state was a pure state. In the present work, we propose an alternative way of defining the quasi inverse using trace distance that allows us to consider mixed input state.
@@
The distance measures between the quantum states play a pivotal role in quantum information science. This is used to compute quantum correlations, entanglement, and coherence as discussed in [6]. The distance measures come into play for several critical tasks in quantum information processing. They help us gauge and distinguish different states of entanglement in bipartite and multipartite systems, especially when the observable set is not closed under products [7]. Among these measures, the trace distance and fidelity are frequently used [1, 8]. However, fidelity is usually easier to calculate only when one of the state is a pure state. On the other hand, the trace distance does not suffer from this drawback. Therefore, to extend the concept of quasi inverse for mixed input states, we propose the mean square trace distance (MSTD) averaged over all possible input states as a meteric to be minimized to find the inverse.
@@
The structure of the paper is as follows: In Section II, we define the quasi inverse based on the mean of the trace distance and the general derivation of quasi inverse for qubit channels is presented in Section III. In Section IV, we derive the quasi inverses of a few qubit channels and concluding remarks are presented in Section V.
@@
Figure 1: The schematic showing the qubit channel $\mathcal{E}$ with mixed input state $\rho$ and its quasi inverse $\mathcal{E}^{i}$.
II Definition of quasi inverse in terms of trace distance
Consider a general single-qubit channel $\mathcal{E}$ that converts a mixed state $\rho$ to $\zeta=\mathcal{E}(\rho)$. Suppose further that the quasi inverse of this channel is $\mathcal{E}^{i}$ that converts $\zeta$ to $\rho^{\prime}=\mathcal{E}^{i}(\zeta)$ as shown schematically in Fig. 1. The density matrices of the states $\rho$ and $\rho^{\prime}$ can be described in terms of the Bloch vectors $\mathbf{r}$ and $\mathbf{r}^{\prime}$, respectively, as $\rho=(I+\mathbf{r}\cdot\mathbf{\sigma})/2$ and $\rho^{\prime}=(I+\mathbf{r}^{\prime}\cdot\mathbf{\sigma})/2$ with $\mathbf{\sigma}$ being the vector of the Pauli operators [1]. Similarly, $\zeta=(I+\mathbf{z}\cdot\mathbf{\sigma})/2$.
We define the mean of the square of the trace distance (MSTD) between $\rho$ and $\rho^{\prime}$ as
$\overline{D^{2}}(\rho,\rho^{\prime})=\frac{1}{4}\int\absolutevalue{\mathbf{r}-\mathbf{r}^{\prime}}^{2}~{}\mathrm{d}^{3}\mathbf{r}\,,$ (1)
@@ II Definition of quasi inverse in terms of trace distance
The channel $\mathcal{E}^{i}$ will be quasi inverse if it minimizes MSTD in Eq. (1) in the optimal way, regardless of the input state. Therefore, similar to the definition of the quasi inverse based on fidelity [2], we define the quasi inverse $\mathcal{E}^{i}$ as a channel that reduces the MSTD between the state $\rho$ and the output state of the channel by at least the same or more than any other channel $\mathcal{E}^{\prime}$, i.e.,
$\overline{D^{2}}(\mathcal{E}^{i}(\zeta),\rho)\leq\overline{D^{2}}(\mathcal{E}^{\prime}(\zeta),\rho)\qquad\forall\quad\mathcal{E}^{\prime}$ (2)
@@
$\mathcal{E}^{i}(\zeta)=V\zeta V^{\dagger}\,,$ (3)
@@
Finding the quasi inverse is essentially finding four real paramters $(x_{0},\mathbf{x})$ that specifies the unitary $V$. To do this, we can setup a constrained optimization problem to maximise $\Delta\overline{D^{2}}$ (the decrease in the MSTD)
$\Delta\overline{D^{2}}=\overline{D^{2}}(\rho,\zeta)-\overline{D^{2}}(\rho,\rho^{\prime})\,$ (5)
over all unitary maps, i.e., maximizing over the real parameters $(x_{0},\mathbf{x})$, subject to the constraint $x_{0}^{2}+\mathbf{x}\cdot\mathbf{x}=1$, for a given channel $\mathcal{E}$.
III General Derivation
When the state passes through an arbitrary channel $\mathcal{E}$, the Bloch vector of the qubit state is transformed as [1]
$\mathbf{r}\longrightarrow\mathbf{z}=M\mathbf{r}+\mathbf{c}\,,$ (6)
where $\mathbf{M}$ is a real $3\times 3$ matrix and $\mathbf{c}$ is a real vector in $\mathbb{R}^{3}$ given by
$M_{ij}=\frac{1}{2}\text{Tr}(\sigma_{i}\mathcal{E}(\sigma_{j}))\,,\qquad c_{i}=\frac{1}{2}\text{Tr}(\sigma_{i}\mathcal{E}(\mathbb{I}))\,.$ (7)
Hence, a quantum channel can be characterized by the pair $(\mathbf{M},\mathbf{c})$, and composition of quantum channel is reflected in the composition of the transformation matrices [1]
@@ III General Derivation
Using Eq. (1), the MSTD of the channel between $\rho$ and $\zeta$ is found as
$\overline{D^{2}}(\rho,\zeta)=\frac{1}{20}\left(\text{Tr}(MM^{\dagger})-2\text{Tr}(M)+3\right)+\frac{1}{4}\absolutevalue{\mathbf{c}}^{2}\,.$ (9)
To evaluate the MSTD integral, the following identities were used:
$\int r_{i}~{}\mathrm{d}V=0\,,\qquad\int r_{i}r_{j}~{}\mathrm{d}V=\frac{1}{5}\delta_{ij}\,,$ (10)
$\begin{gathered}\mathrm{d}V=r^{2}\sin\theta~{}\mathrm{d}r~{}\mathrm{d}\theta~{}\mathrm{d}\phi\,,\\ \mathbf{r}=\langle r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta\rangle\,,\\ r\in[0,1]\,,\quad\theta\in[0,\pi]\,,\quad\phi\in[0,2\pi]\,.\end{gathered}$ (11)
The MSTD of combined channel $\mathcal{E}^{i}\circ\mathcal{E}$ is given by
$\overline{D^{2}}(\rho,\rho^{\prime})=\frac{1}{20}\left(\text{Tr}(NN^{\dagger})-2\text{Tr}(N)+3\right)+\frac{1}{4}\absolutevalue{\mathbf{u}}^{2}\,,$ (12)
where $N=M^{i}M$ and $\mathbf{u}=M^{i}\mathbf{c}+\mathbf{c}^{i}$.
@@
IV Examples
To find the qusi inverse of specific channels, we consider the same example channels as in Ref. [2] for comparison of the results.
IV.1 The Pauli Channel

### v3

@@ Abstract
We found the quasi inverse of qubit channels as a unitary map, $\mathcal{E}^{i}$, by minimizing the average trace distance between the input state to the channel and the output of the quasi inverse channel for arbitrary qubit channel $\mathcal{E}$ and for arbitrary input states. The channel $\mathcal{E}$ was assumed completely positive and trace-preserving. To find the quasi inverse for mixed states, we proposed an alternative definition of the quasi inverse based on the mean square of the trace distance (MSTD) of a channel. The definition based on the trace distance allowed easy generalization of the quasi inverse to mixed input states. The quasi inverse of the Pauli, generalized amplitude damping, mixed unitary, and tetrahedron channels calculated based on the MSTD agreed with the one computed using average fidelity in the special case of input states being pure.
@@ I Introduction
Qubits are the building blocks of quantum computers, but they are quite vulnerable to unwanted interactions with the environment. This interaction leads to the decoherence of the quantum state and makes the computations erroneous on quantum computers. This interaction process can be modeled as a quantum channel [1]. In general, quantum channels are irreversible operations [2, 3, 4]. However, the idea of quasi inversion of the quantum channel was recently introduced [2] to reverse a part of the effect of that channel. The concept of quasi inverse is essential for enhancing the fidelity of quantum gates and reducing errors in quantum computations. This idea is also helpful in quantum inverse problems, such as state estimation, where quasi inverses are used to improve efficiency and resilience [5].
@@
The quasi inverse is defined so as to increase the average fidelity of the channel in Ref. [2]. The derivation of the quasi inverse was based on maximizing the average fidelity between the input state of the channel and the output state of the quasi inverse of that channel. However, the derivation was restricted to the case when the input state was a pure state. In the present work, we propose an alternative way of defining the quasi inverse using the trace distance that allows us to consider mixed input state.
@@
The distance measures between the quantum states play a pivotal role in quantum information science. This is used to compute quantum correlations, entanglement, and coherence as discussed in [6]. The distance measures come into play for several critical tasks in quantum information processing. They help us gauge and distinguish different states of entanglement in bipartite and multipartite systems, especially when the observable set is not closed under products [7]. Among these measures, the trace distance and fidelity are frequently used [1, 8]. However, fidelity is usually easier to calculate only when one of the state is a pure state. On the other hand, the trace distance does not suffer from this drawback. Therefore, to extend the concept of quasi inverse for mixed input states, we propose the mean square trace distance (MSTD): the square of the trace distance averaged over all possible input states.
@@
The structure of the paper is as follows: In Section II, we construct the MSTD for a channel from the definition of the trace distance between two density matrices. In Section III, we define the quasi inverse using the MSTD and the general derivation of quasi inverse for qubit channels is presented in Section IV. In Section V, we derive the quasi inverses of a few qubit channels and concluding remarks are presented in Section VI.
II Mean square trace distance (MSTD)
Consider two single-qubit mixed states $\rho$ and $\zeta$. The density matrices $\rho$ and $\zeta$ can be represented as geometric vectors in the unit Bloch ball with their position vectors $\mathbf{r}$ and $\mathbf{z}$ having components $r_{i}={\rm Tr}(\rho\sigma_{i})$ and $z_{i}={\rm Tr}(\zeta\sigma_{i})$, $i=1,2,3,$ where $\sigma_{i}$ are the Pauli operators, $0\leq r\leq 1$, and $0\leq z\leq 1$. Therefore, we can express the density matrices in terms of the Bloch vectors as
$\rho=\frac{1}{2}(I+\mathbf{r}\cdot\mathbf{\sigma})$
and
$\zeta=\frac{1}{2}(I+\mathbf{z}\cdot\mathbf{\sigma})$
with $\mathbf{\sigma}$ being the vector of the Pauli operators and $I$ being the identity operator.
The trace distance between two single-qubit mixed states $\rho$ and $\zeta$ is defined as [1]
$D(\rho,\zeta)$ $=$ $\frac{1}{2}{\rm Tr}\,\absolutevalue{\rho-\zeta}$ | (1)
$=$ $\frac{1}{4}{\rm Tr}\absolutevalue{(\mathbf{r}-\mathbf{z})\cdot\mathbf{\sigma}}$ |
$=$ $\frac{1}{2}\absolutevalue{\mathbf{r}-\mathbf{z}}$ |
where we used the fact that the eigenvalues of $(\mathbf{r}-\mathbf{z})\cdot\sigma$ are ±|$\pm\absolutevalue{\mathbf{r}-\mathbf{z}}$ |.
Now consider a qubit channel $\mathcal{E}$ as shown schematically in Fig. 1. The MSTD of the channel is defined as
$\overline{D^{2}}(\mathcal{E})=\int D^{2}\left(\rho,\mathcal{E}(\rho)\right)\,\mathrm{d}\rho\,,$ (2)
where the integration is to be performed over all single-qubit mixed states.
@@ II Mean square trace distance (MSTD)
Figure 1: The schematic showing the qubit channel $\mathcal{E}$ with mixed input state $\rho$ and output state $\mathcal{E}(\rho)$.
The quantum channel $\mathcal{E}$ can be represented as an affine transformation of the Bloch vectors of the input density matrices to the Bloch vectors of the output density matrices as [1]
$\mathbf{r}\longrightarrow\mathbf{z}=M\mathbf{r}+\mathbf{c}\,,$ (3)
where ${M}$ is a real $3\times 3$ matrix and $\mathbf{c}$ is a real vector in $\mathbb{R}^{3}$ given by
$M_{ij}$ $=$ $\frac{1}{2}\text{Tr}(\sigma_{i}\mathcal{E}(\sigma_{j}))\,,\quad i,j=1,2,3,$ (4)
$c_{i}$ $=$ $\frac{1}{2}\text{Tr}(\sigma_{i}\mathcal{E}(I))\,,\quad i=1,2,3\,.$ (5)
Hence, a quantum channel can be characterized by the pair $({M},\mathbf{c})$. Therefore, using Eq. (1), the definition (2) can be recast as
$\overline{D^{2}}(\mathcal{E})$ $=$ $\frac{1}{4}\int\absolutevalue{\mathbf{r}-\mathbf{z}}^{2}~{}\mathrm{d}^{3}\mathbf{r}\,,$ (6)
$=$ $\frac{1}{4}\int\absolutevalue{\mathbf{r}-\left(M\mathbf{r}+\mathbf{c}\right)}^{2}~{}\mathrm{d}^{3}\mathbf{r}\,,$ (7)
@@
III Quasi inverse in terms of MSTD
Consider a general single-qubit channel $\mathcal{E}$ that converts a mixed state $\rho$ to $\mathcal{E}(\rho)$. Suppose further that another channel $\mathcal{E}^{\prime}$ transforms $\mathcal{E}(\rho)$ to $\rho^{\prime}=(I+\mathbf{r}^{\prime}\cdot\mathbf{\sigma})/2=\mathcal{E}^{\prime}\circ\mathcal{E}(\rho)$ such that
$\overline{D^{2}}\big{(}\mathcal{E}^{\prime}\circ\mathcal{E}\big{)}\leq\overline{D^{2}}\big{(}\mathcal{E}\big{)}\,.$ (8)
as shown schematically in Fig. 2. The quasi inverse $\mathcal{E}^{i}$ is a channel such that
$\mathcal{E}^{i}=\max\limits_{\Delta\overline{D^{2}}}\,\,\mathcal{E}^{\prime}$ (9)
where
$\Delta\overline{D^{2}}=\overline{D^{2}}\big{(}\mathcal{E}\big{)}-\overline{D^{2}}\big{(}\mathcal{E}^{\prime}\circ\mathcal{E}\big{)}\,.$ (10)
[…]
Figure 2: The schematic showing quasi inverse of the qubit channel $\mathcal{E}$.
Therefore, the channel $\mathcal{E}^{\prime}$ will be quasi inverse if it maximizes $\Delta\overline{D^{2}}(\mathcal{E}^{\prime}\circ\mathcal{E})$ for a given channel $\mathcal{E}$. Let us note that it is equivalent to either minimize $\overline{D^{2}}(\mathcal{E}^{\prime}\circ\mathcal{E})$ or maximize $\Delta\overline{D^{2}}$ over all CPTP maps $\mathcal{E}^{i}$. However, maximizing $\Delta\overline{D^{2}}$ is preferred over minimizing $\overline{D^{2}}(\mathcal{E}^{\prime}\circ\mathcal{E})$, as both yield equivalent quasi inverses, but the former produces equations with a clear and interpretable structure, consistent with Ref. [2] and the calculations are simpler. Therefore, similar to the definition of the quasi inverse based on fidelity [2], we define the quasi inverse $\mathcal{E}^{i}$ as a channel that reduces the MSTD between its output state and the input state of the original channel $\mathcal{E}$ by at least the same or more than any other channel $\mathcal{E}^{\prime}$, i.e.,
$\overline{D^{2}}\big{(}\mathcal{E}^{i}\circ\mathcal{E}\big{)}\leq\overline{D^{2}}\big{(}\mathcal{E}^{\prime}\circ\mathcal{E}\big{)}\qquad\forall\qquad\mathcal{E}^{\prime}\,.$ (11)
@@ IV General Derivation
$\mathcal{E}^{i}(\rho)=V\rho V^{\dagger}\,,$ (12)
@@
Finding the quasi inverse is essentially finding four real parameters $(x_{0},\mathbf{x})$ that specifies the unitary $V$. To do this, we can set up a constrained optimization problem to maximize $\Delta\overline{D^{2}}$ (the decrease in the MSTD) over all unitary maps, i.e., maximizing over the real parameters $(x_{0},\mathbf{x})$, subject to the constraint $x_{0}^{2}+\mathbf{x}\cdot\mathbf{x}=1$, for a given channel $\mathcal{E}$.
Using Eq. (6), the MSTD of the channel $\mathcal{E}$ is found as
$\overline{D^{2}}(\mathcal{E})=\frac{1}{20}\left(\text{Tr}(MM^{\dagger})-2\text{Tr}(M)+3\right)+\frac{1}{4}\absolutevalue{\mathbf{c}}^{2}\,.$ (14)
To evaluate the MSTD integral, the following identities were used:
$\int r_{i}~{}\mathrm{d}^{3}\mathbf{r}=0\,,\qquad\int r_{i}r_{j}~{}\mathrm{d}^{3}\mathbf{r}=\frac{1}{5}\delta_{ij}\,,$ (15)
$\begin{gathered}\mathrm{d}^{3}\mathbf{r}=r^{2}\sin\theta~{}\mathrm{d}r~{}\mathrm{d}\theta~{}\mathrm{d}\phi\,,\\ \mathbf{r}=(r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta)\,,\\ r\in[0,1]\,,\quad\theta\in[0,\pi]\,,\quad\phi\in[0,2\pi]\,.\end{gathered}$ (16)
The MSTD of combined channel $\mathcal{E}^{i}\circ\mathcal{E}$ is given by
$\overline{D^{2}}\big{(}\mathcal{E}^{i}\circ\mathcal{E}\big{)}=\frac{1}{20}\left(\text{Tr}(NN^{\dagger})-2\text{Tr}(N)+3\right)+\frac{1}{4}\absolutevalue{\mathbf{u}}^{2}\,,$ (17)
where
$N$ $=M^{i}M$ (18)
$\mathbf{u}$ $=M^{i}\mathbf{c}+\mathbf{c}^{i}$ (19)
since the composition of a quantum channel is reflected in the composition of the transformation matrices [1]
@@
The affine map for the quasi inverse channel (12) can be written using simple algebra as
$\begin{gathered}M^{i}=\begin{pmatrix}1-2(x_{2}^{2}+x_{3}^{2})&2(x_{0}x_{3}+x_{1}x_{2})&-2(x_{0}x_{2}-x_{1}x_{3})\\ -2(x_{0}x_{3}-x_{1}x_{2})&1-2(x_{1}^{2}+x_{3}^{2})&2(x_{0}x_{1}+x_{2}x_{3})\\ 2(x_{0}x_{2}+x_{1}x_{3})&-2(x_{0}x_{1}-x_{2}x_{3})&1-2(x_{1}^{2}+x_{2}^{2})\end{pmatrix}\,,\\ \mathbf{c}^{i}=\begin{pmatrix}0&0&0\end{pmatrix}^{T}\,.\end{gathered}$ (21)
The last equality is due to the fact that the quasi inverse is taken to be a unitary channel.
@@
V Examples
To find the quasi inverse of specific channels, we consider the same example channels as in Ref. [2] for comparison of the results. However, we begin with a sanity check and find the quasi inverse of a general unitary channel since its inverse can be exactly written using the adjoint of unitary operator defining the channel. This will also illustrate the process of finding quasi inverse clearly.
V.1 General Unitary Operator
A general unitary operator is a rotation of the qubit state by some angle $\theta$ about an arbitrary axis $\hat{n}$ and can be written as
$U=e^{-i\theta\mathbf{n}\cdot\sigma/2}=\cos\left(\frac{\theta}{2}\right)I-i\sin\left(\frac{\theta}{2}\right)\left[n_{1}\sigma_{1}+n_{2}\sigma_{2}+n_{3}\sigma_{3}\right]$ (22)
where $\mathbf{n}=\left(n_{1},n_{2},n_{3}\right)$ is a real unit vector in $\mathbb{R}^{3}$. The general unitary channel defined as
$\mathcal{E}(\rho)=U\rho U^{\dagger}$ (23)
can also be equivalently defined using the affine map like that in Eq. (21) as
$M=\begin{pmatrix}\beta_{x}&\alpha_{xyz}&\alpha_{xzy}&\\ \alpha_{xyz}&\beta_{y}&\alpha_{xzy}\\ \alpha_{xzy}&\alpha_{yzx}&\beta_{z}\end{pmatrix}$ (24)
where $\mathbf{c}=(0,0,0)^{T}$, $\beta_{i}=2n_{i}^{2}\sin^{2}\left(\frac{\theta}{2}\right)+\cos(\theta)$ and $\alpha_{ijk}=2n_{i}n_{j}\sin^{2}\left(\frac{\theta}{2}\right)-n_{k}\sin(\theta)$. Using Eqs. (14), (17), (21), and (24) in Eq. (10) we get
$\begin{gathered}\Delta\overline{D^{2}}=\frac{2}{5}\left[\omega_{x}x_{1}^{2}+\omega_{y}x_{2}^{2}+\omega_{z}x_{3}^{2}\right.\\ +2\epsilon_{x}x_{0}x_{1}+2\epsilon_{y}x_{0}x_{2}+2\epsilon_{z}x_{0}x_{3}\\ \left.2\mu_{xy}x_{1}x_{2}+2\mu_{xz}x_{1}x_{3}+2\mu_{yz}x_{2}x_{3}\right]\,,\end{gathered}$ (25)
where $\epsilon_{i}=\frac{1}{2}n_{i}\sin(\theta)$, $\omega_{i}=n_{i}^{2}\sin^{2}\left(\frac{\theta}{2}\right)+\sin^{2}\left(\frac{\theta}{2}\right)-1$, and $\mu_{ij}=n_{i}n_{j}\sin^{2}\left(\frac{\theta}{2}\right)$. This allows us to write the above expression in matrix representation given by:
where
$Q=\begin{pmatrix}0&\epsilon_{x}&\epsilon_{y}&\epsilon_{z}\\ \epsilon_{x}&\omega_{x}&\mu_{xy}&\mu_{xz}\\ \epsilon_{y}&\mu_{xy}&\omega_{y}&\mu_{yz}\\ \epsilon_{z}&\mu_{xz}&\mu_{yz}&\omega_{z}\end{pmatrix}\,.$ (27)
The eigenvalues of $Q$ are given by:
$\lambda_{1}=\lambda_{2}=\lambda_{3}$ $=-\cos^{2}\left(\frac{\theta}{2}\right),\quad\lambda_{4}=\sin^{2}\left(\frac{\theta}{2}\right)$ (28)
Thus, the decrease in the MSTD is given by:
$\Delta\overline{D^{2}}$ $=\frac{2}{5}\max\left(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\right)$ (29)
$=\frac{2}{5}\sin^{2}\left(\frac{\theta}{2}\right)$
$=D^{2}(\mathcal{E})$
The eigenstate corresponding to the largest eigenvalue is our solution as explained in the Appendix A. The largest eigenvalue $\lambda_{\max}=\lambda_{4}$ corresponds to the eigenstate
|$\ket{\lambda_{\max}}=\begin{pmatrix}\cos\left(\frac{\theta}{2}\right)\\[5.69054pt] n_{x}\sin\left(\frac{\theta}{2}\right)\\[5.69054pt] n_{y}\sin\left(\frac{\theta}{2}\right)\\[5.69054pt] n_{z}\sin\left(\frac{\theta}{2}\right)\end{pmatrix}\,.$ (30)
Thus the quasi inverse is:
$V$ $=\cos\left(\frac{\theta}{2}\right)I+i\sin\left(\frac{\theta}{2}\right)\left(\mathbf{n}\cdot\mathbf{\sigma}\right)$ (31)
$=e^{i\theta\mathbf{n}\cdot\mathbf{\sigma}}$
$=U^{\dagger}$
after using the fact that $(x_{0},\mathbf{x})=(x_{0},x_{1},x_{2},x_{3})=\ket{\lambda_{\max}}$ and Eq. (13).
V.2 The Pauli Channel